# Two-Player Zero-Sum Games with Bandit Feedback

### v1

@@ Appendix A Subgaussian Properties
We note that $\Delta_{ij}^{*}\leq\Delta_{ij}$, which ensures that the regret does not grow excessively, and this relationship allows us to analyze the regret in terms of the defined action sets. As a result, if we sum these regret contributions as mentioned, we can conclude the proof.
$\Box$
C.1 Proof of Theorem 5.2
We consider the regret approach consisting of the losses of both players by comparing their best choices to the action played. This regret measures how much worse a player performs compared to their optimal strategy if they had known the opponent’s behavior in advance. Clearly we can write the followings:
$R_{T}$ $=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{max}}\mathbb{E}[n_{ij,T}]+\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{min}}\mathbb{E}[n_{ij,T}]$
$=\sum_{(i,j)\in S_{A}}\Delta_{ij}\mathbb{E}[n_{ij,T}]$
$=\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}\mathbb{E}[n_{ij,T}]+\sum_{(i,j)\in S_{A_{2}}}\Delta_{ij}\mathbb{E}[n_{ij,T}]$
This implies that the term involving $\Delta_{ij}^{\text{max}}$ characterizes the regret incurred by the maximizing player, whereas the loss of the minimizing player is determined using $\Delta_{ij}^{\text{min}}$. In particular, we utilize the same underlying algorithm, which maintains the same uniform playing strategy and elimination strategy throughout the process. This consistency allows us to to leverage many elements of the previous regret analysis. Consequently, we can apply the same case scenarios from the proof of Theorem 5.1 as the elimination of action pairs follows the same strategy and the selection process of action pairs remains the same across rounds.
Although our current analysis involves a different notion of regret, consisting of $\Delta_{ij}$, this does not affect the underlying structure of the regret analysis. This ensures that the previous theoretical bounds can be extended by using $\Delta_{ij}$ instead of $\Delta_{ij}^{*}$ and we apply $\Delta_{ij}^{*}\leq\Delta_{ij}$ and $\Delta_{ij}\leq\lambda$ for the action pairs in $S_{A_{2}}$, then the result follows.
$\Box$
Appendix D Experimental Details
@@
In Figure 1(b), we compare the performance of the ETC-TPZSG, ETC-TPZSG-AE and Tsallis-INF [18] algorithms. In each run, the exploration time $k$ for each action pair in ETC-TPZSG is randomly selected from a predefined list ranging from 100 to 2500 in increments of 100. The total number of rounds is set to $T=10^{4}$. On the other hand, each run uses a new game matrix with a unique pure NE which is generated using Gaussian payoffs with mean zero, standard deviation $\sigma$ selected from the list $[0.25,0.5,0.75,1]$. Varying $\sigma$ not only changes the payoff distribution but also affects the exploration time and the tolerance parameter $\varepsilon$ used in the elimination phase of ETC-TPZSG-AE. The cumulative regret is averaged over $10^{3}$ simulation runs. As observed, the ETC-TPZSG-AE algorithm consistently outperforms ETC-TPZSG, demonstrating the effectiveness of its elimination strategy in reducing cumulative regret, while Tsallis-INF perform better when compared to ETC-based algorithms. Notably, the ETC-based algorithms offer the advantage of conceptual and implementational simplicity.
Figure 2: Theoretical expected regret bound comparison between ETC-TPZSG, ETC-TPZSG-AE and Tsallis-INF from [18] using different $\Delta$ values
On the other hand, Figure 2 compares the theoretical expected regret bounds (in $O(\cdot)$ notation) of the algorithms across different values of the suboptimality gap $\Delta$. The results show that Tsallis-INF tends to perform better when $\Delta$ is large while the ETC-TPZSG-AE algorithm achieves lower regret for smaller values of $\Delta$.

### v2

@@ Appendix C Proof of Theorem 2
We note that $\Delta_{ij}^{*}\leq\Delta_{ij}$, which ensures that the regret does not grow excessively, and this relationship allows us to analyze the regret in terms of the defined action sets. As a result, if we sum these regret contributions as mentioned, we can conclude the proof. $\Box$
Appendix D Proof of Theorem 4
We follow exactly the same steps as the proof of Theorem 2. We update the threshold parameter as $\lambda\geq 4\sigma\sqrt[4]{e/T}$, thus we have two subsets of action pairs such that $S_{A_{1}}=\{(i,j)\in S_{A}:\Delta_{ij}>\lambda\}$ and $S_{A_{2}}=\{(i,j)\in S_{A}:0<\Delta_{ij}\leq\lambda\}$. We can use the same probability of keeping an action pair as in (57) since the algorithm each action pair at least $k_{t}$ times. Then, if we keep the same events from the proof of Theorem 2, Appendix C, we have the following:
$\leq\exp{\Bigg(-\frac{\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}{2(\frac{\sigma^{2}}{k_{i^{\prime}j,t}}+\frac{\sigma^{2}}{k_{ij,t}})}\Bigg)}$ (68)
$\leq\exp{\Bigg(-\frac{\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}{2\Big(\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{i^{\prime}j,t})}+\frac{\sigma^{2}}{(k_{t}+k_{t}\hat{P}_{ij,t})}\Big)}\Bigg)}$ (69)
$\leq\exp{\Bigg(-\frac{\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}{2\sigma^{2}\Big(\frac{(\hat{P}_{i^{\prime}j,t}+\hat{P}_{ij,t}+2)}{k_{t}(1+\hat{P}_{i^{\prime}j,t})(1+\hat{P}_{ij,t})}\Big)}\Bigg)}$ (70)
$=\exp{\Bigg(-\frac{2(1+\hat{P}_{i^{\prime}j,t})(1+\hat{P}_{ij,t})}{(\hat{P}_{i^{\prime}j,t}+\hat{P}_{ij,t}+2)}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}\Bigg)}$ (71)
$\leq\exp{\Bigg(-\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}\Bigg)$ (72)
$=\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}T}$ (73)
Since the term $\frac{2(1+\hat{P}_{i^{\prime}j,t})(1+\hat{P}_{ij,t})}{(\hat{P}_{i^{\prime}j,t}+\hat{P}_{ij,t}+2)}$ is greater than one, the inequality (72) follows. In addition, since we obtain the same result for $\mathbb{P}(E_{2})$ as $\mathbb{P}(E_{2})$, which leads to the same probability of keeping an action pair as in the proof of Theorem 2 and also the same results for Case 1, Case 3.
On the other hand, since we explore each action pair non-uniformly, we need to update Case 2 by using
$k_{t_{ij}}=\Big\lceil\frac{16\sigma^{2}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil+\Big\lceil\frac{16\sigma^{2}\hat{P}_{ij,t_{ij}}}{\hat{\Delta}_{t_{ij}}^{2}}\ln{(\frac{\hat{\Delta}_{t_{ij}}^{2}T}{16\sigma^{2}})}\Big\rceil\leq 2\ \Big\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big\rceil$
since $\hat{P}_{ij,t_{ij}}\leq 1$ for all $i,j$. Then, the regret contribution becomes
$\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\ 2\ \Big\lceil\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big\rceil$ $\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}+\Delta_{ij}^{*}\frac{512\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}$ (74)
Then, it concludes the proof.
Theorem 6.
The external regret of Algorithm 3 is bounded by
$R_{T}\leq$ $\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}+\frac{768\sigma^{2}}{\Delta_{ij}}+\frac{512\sigma^{2}}{\Delta_{ij}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\frac{512\sigma^{2}}{\lambda}+\lambda T\Big).$
Since we have $\Delta_{ij}^{*}\leq\Delta_{ij}$, it concludes the proof.
Appendix E Experimental Details
@@
In Figures 3 and 3 , we compare the performance of the ETC-TPZSG, ETC-TPZSG-AE, ETC-TPZSG-AE-NUE and Tsallis-INF [18] algorithms. In each run, the exploration time $k$ for each action pair in ETC-TPZSG is randomly selected from a predefined list ranging from 100 to 2500 in increments of 25. The total number of rounds is set to $T=10^{4}$. On the other hand, each run uses a new game matrix with a unique pure NE which is generated using Gaussian payoffs with mean zero, standard deviation $\sigma$ selected from the list $[0.1,0.2]$ for Figure 3 and from the list $[0.25,0.5,0.75,1]$ for Figure 3. Here, $\sigma$ denotes the variance factor of Gaussian payoffs and setting it to a small value leads to smaller $\Delta$ values. Furthermore, varying $\sigma$ not only changes the payoff distribution but also affects the exploration time and the tolerance parameter $\varepsilon$ used in the elimination phase of ETC-TPZSG-AE. The cumulative regret is averaged over $10^{2}$ simulation runs. Furthermore, Figures 4 and 5 present the comparison of algorithms based on different game matrices. For large $\Delta$, although the Tsallis-INF algorithm performs better in the first $10^{3}$ rounds, over time, our adaptive elimination algorithms reach lower cumulative regret.
As observed, the ETC-TPZSG-AE algorithm consistently outperforms ETC-TPZSG, demonstrating the effectiveness of its elimination strategy in reducing cumulative regret, and ETC-TPZSG-AE-NUE outperforms even for larger action sets. Increasing the number of actions for players makes the identification of the true NE more challenging, as reflected in the performance of ETC-TPZSG as shown in Figures 4 and 5. On the other hand, when $\Delta$ is smaller, Tsallis-INF [18] also has difficulty in identifying pure NE correctly, as seen in Figure 5. However, Algorithms 2 and 3 perform well even in these cases, as they incorporate pure NE properties during the elimination phase. In particular, ETC-based algorithms offer the advantage of conceptual and implementational simplicity.
Figure 4: The cumulative regrets of the algorithms for different game matrices with large $\Delta$
Figure 5: The cumulative regrets of the algorithms for different game matrices with small $\Delta$